OpenStudy (anonymous):

Which equation is a quadratic equation? y + 7 = −3x(4 − x) + 9x y = −2x + 5 y + 2x3 = (2x − 2)(x + 5) y − 9x = (x2 + 1)(5x − 1) + 3

OpenStudy (anonymous):

y+7=-3x(4-x)+9x y+7=-12x+3x^2+9x y+7=3x^2-3x y=3x^2-3x-7

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

So do you get it?

OpenStudy (anonymous):

do it would be answer d

OpenStudy (anonymous):

Remember that the quadratic equation is ax^2+bx+c=0

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

No problem